A supermarket boasts that checkout times for customers are never more than 15 minutes. A random sample of 36 costumers reveals a mean checkout time of 17 minutes with a standard deviation of 3 minutes. What can you conclude about the supermarket’s boast at the 0.05 level? 


1
Expert's answer
2022-05-05T17:01:39-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le15"

"H_a:\mu>15"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=35" degrees of freedom, and the critical value for a right-tailed test is "t_c = 1.689572."

The rejection region for this right-tailed test is "R = \{t: t > 1.689572\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{17-15}{3/\sqrt{36}}=4"

Since it is observed that "t = 4 >1.689572= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for right-tailed, "df=35" degrees of freedom, "t=4" is "p = 0.000156," and since "p= 0.000156<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 15, at the "\alpha = 0.05" significance level.